How to solve these ? 


1)If the equation(a^2+b^2)x^2 - 2b(a+c)x + b^2+c^2 has equal roots then , which of the following is right and how ?


    a) 2b=a+c b)b^2=ac c)b=2ac/a+c d)b=ac


2)If the equation(a^2+b^2)x^2 - 2(ac+bd)x +c^2+d^2 = 0 has equal roots then


    a)ab=cd b)ad=bc c)ad=under root bc d)ab=under root cd


Grade:10

2 Answers

Mostafijur Rahaman
10 years ago
1)b)b^2=ac 2)b)ad=bc
Pushkar Aditya
10 years ago
1)as it has equal roots so,(2ba+2cb)^2=4(a^2+b^2)(b^2+c^2) 8ab^2c=4b^2(b^2+c^2) (b-c)^2=0 b=c 2ac=2b^2 b)b^2=ac answer 2)(2ac+2bd)^2=4(a^2+b^2)(c^2+d^2) (ad-bc)^2=0 b)ad=bc answer